Find the relative extremes of f(x,y)= x - x2y - xy2
Answer #1

2 & fy: 0 to = 1-2xy-y² 2=0 And flory) = x - xy - xy For Cntical points. Put te=0 ty = -x ² - 2xy Now. Hy so 1-2xy = y² = 0 2

Please appreciate my work.
If you have any query regarding this question. You are free to ask.